OpenStudy (anonymous):

simplify the equation 3y-5x=15 and how do i graph it

OpenStudy (anonymous):

Write y as a function of x and then vary x as independent variable and get you plot these points on cartesian

OpenStudy (anonymous):

Put it into y=mx+b form. First get "y" by itself. You should end up with Y=5/3X+3. +3 is your y-intercept and the slope is (rise over run) 5/3

OpenStudy (anonymous):

graph using x and y intercept

OpenStudy (anonymous):

what is x intercept

OpenStudy (anonymous):

Just plug "0" in for x and y. 3(0)-5x=15 solve for x. and then 3y-5(0)=15 solve for y. 3x0=0, -5x=15 (divide by -5) x=-3 so your first point is (-3,0) -5x0=0, 3y=15 (divide by 3) y=3 so your next point is (0,3)